OpenStudy (anonymous):

Translate the sentence into an equation. The difference of a number times 2 and 9 is equal to 3 . Use the variable x for the unknown number.

OpenStudy (anonymous):

\[\left| 2x -9 \right| = 3\]

OpenStudy (anonymous):

Don't forget to acknowledge w/a medal

OpenStudy (anonymous):

X x 2 + 9 = 3

OpenStudy (anonymous):

wait sorry + is supossed to be -

OpenStudy (anonymous):

the mod has to be there because it is given difference b/w the two and it is not specified which one would be greater 2x or 9